# Vision Transformers in Precision Agriculture: A Comprehensive Survey

### v1

@@ 4.1 Pure Models
A simple use-case of ViT is shown in the study by [52], which proposed a ViT-based approach for the automatic classification of strawberry diseases using transfer learning and fine-tuning techniques. The authors enhanced the ViT model by adding new layers such as ReLU activation, batch normalization, dropout, and a softmax classifier. Comparative evaluations against conventional CNN models like VGG16, VGG19, ResNet50V2, and MobileNet showed that the proposed ViT model, especially when combined with data augmentation, significantly outperformed other methods. In their best setting, the model achieved an accuracy and F1-score of 92.7% on the test set. The research highlights the strengths of Transformer-based architectures in precision agriculture, especially in complex visual tasks such as disease recognition from strawberry leaves, fruits, and flowers. The study also evaluated the proposed method on the PlantVillage dataset to assess generalizability, where it attained 98.9% accuracy, outperforming both the original ViT and previous state-of-the-art methods like Mask R-CNN and MCLCNN. The authors concluded that the ViT model’s ability to capture global image features makes it a powerful alternative to CNNs for agricultural disease detection tasks
@@
In addition to using a custom dataset, [99] proposed a novel residual-distilled Transformer architecture for the identification of rice leaf diseases. The authors argue that existing works on rice leaf disease identification are unsatisfactory, either in terms of identification accuracy or model interpretability. To address these issues, they introduce a new method that combines a ViT and a distilled Transformer in a residual block for feature extraction, followed by a MLP for prediction. The researchers collected their own dataset on rice leaf disease from the years 2019 to 2020. This dataset, which they refer to as the rice leaf disease dataset, consists of pictures of four rice leaf diseases: bacterial blight, brown spots, blast, and tungro. The proposed method achieved an 89% F1-score and 92% top-1 accuracy on the rice leaf disease dataset, outperforming other state-of-the-art models.
@@
Table 1: Summary of pure models. These models employ the pure form of ViT, aiming to enhance its performance or fine-tune it. (Reported metrics are directly extracted from the original publications. In certain studies, per-class or metrics are provided; metrics that are not reported are indicated with a "-".)
@@ 4.2 Hybrid Models
[41] Self-built 99.10% 0.9777 0.9847 0.9775 IUsing Multi-Head Self-Attention (MSA) and a Spatial Convolutional Self-Attention-based Transformer (SCSA-Transformer) for accurate and efficient recognition of multiple strawberry disease classes
@@ 5.1 Key Findings
  • Self-attention mechanism is a key advantage: The self-attention mechanism allows ViTs to focus on the most important parts of an image without manual feature extraction. This is helpful in detecting small, subtle spots on leaves. ViTs can dynamically weigh different regions of the input image, ensuring that significant features receive adequate attention during classification. This helps capture both local and global contextual information. This approach is different from CNNs s, which apply convolution operations uniformly across an image and can sometimes miss small but significant features. The ability of ViTs to focus on key features helps improve accuracy and reduce the need for extensive pre-processing of images.
@@ 5.2 Open Challenges
Despite the significant progress made with ViTs in precision agriculture, several challenges still hinder their wider adoption. A key issue is the absence of inductive bias in ViTs. Unlike CNNs, which take advantage of spatial hierarchies and local patterns in images through convolutional filters, ViTs do not possess these built-in inductive biases. Consequently, ViTs require large, high-quality datasets to perform well, making them data-hungry. This reliance on vast amounts of data can be a major obstacle in agriculture, where acquiring labeled datasets is both costly and time-consuming. As a result, ViTs may struggle to reach optimal performance without substantial data augmentation or pre-training on large-scale datasets like ImageNet.
@@
  • Need for more comprehensive and diverse datasets: Many studies use publicly available datasets such as PlantVillage, which may not fully represent the variability of real-world scenarios. Real-world conditions involve a multitude of factors, such as diverse plant species, various disease types and stages, and different environmental and lighting conditions—factors that are often not captured in standard datasets. Datasets also need to account for image variations, including different angles, distances, and image quality. The lack of diverse data can lead to models that are not robust and may not generalize well to new, unseen data. Although some studies have created their own datasets to address this issue, collecting and annotating large-scale, high-quality datasets remains expensive and time-consuming. The number of times each dataset has been used is shown in Table LABEL:tab:tabledata.
@@
Table 3: Frequency of dataset usage across the analyzed studies. Datasets created by gathering images from the internet or by combining other datasets are categorized under the ’Custom’ category. Datasets that don’t have a unique name or doi link and are simply accessible from Kaggle are categorized under the ’From Kaggle’ category.
@@
Self-built Dataset that built by the authors $12$
@@
Custom Datasets that are gathered from search engines, created by combining other datasets, or are not easily accessible $8$
@@
CD&S [3] Contains over 4,000 Corn images comprising of field images and augmented images $2$
@@
DRLI [15] Contains over 4000 images of healthy and diseased leafs $1$
PlantComposite [56] Contains over 58,000 images of helathy and disease leaves $1$
Citrus Fruits and Leaves dataset [62] Contains over 600 images of citrus leaves and 150 images of citrus fruits and five classes for each one $1$
Dhan-Shomadhan [31] Contains over 1100 images of five rice leaves diseases $1$
Strawberry Disease dataset [1] includes 2500 images of seven kinds of strawberry diseases $1$

### v2

@@ 4.1 Pure Models
A simple use-case of ViT is shown in the study by [52], which proposed a ViT-based approach for the automatic classification of strawberry diseases using transfer learning and fine-tuning techniques. The authors enhanced the ViT model by adding new layers such as ReLU activation, batch normalization, dropout, and a softmax classifier. Comparative evaluations against conventional CNN models like VGG16, VGG19, ResNet50V2, and MobileNet showed that the proposed ViT model, especially when combined with data augmentation, significantly outperformed other methods. In their best setting, the model achieved an accuracy and F1-score of 92.7% on the test set. The research highlights the strengths of Transformer-based architectures in precision agriculture, especially in complex visual tasks such as disease recognition from strawberry leaves, fruits, and flowers. The study also evaluated the proposed method on the PlantVillage dataset to assess generalizability, where it attained 98.9% accuracy, outperforming both the original ViT and previous state-of-the-art methods like Mask R-CNN and MCLCNN. The authors concluded that the ViT model’s ability to capture global image features makes it a powerful alternative to CNNs for agricultural disease detection tasks.
@@
In addition to using a self-built dataset, [99] proposed a novel residual-distilled Transformer architecture for the identification of rice leaf diseases. The authors argue that existing works on rice leaf disease identification are unsatisfactory, either in terms of identification accuracy or model interpretability. To address these issues, they introduce a new method that combines a ViT and a distilled Transformer in a residual block for feature extraction, followed by a MLP for prediction. The researchers collected their own dataset on rice leaf disease from the years 2019 to 2020. This dataset, which they refer to as the rice leaf disease dataset, consists of pictures of four rice leaf diseases: bacterial blight, brown spots, blast, and tungro. The proposed method achieved an 89% F1-score and 92% top-1 accuracy on the rice leaf disease dataset, outperforming other state-of-the-art models.
@@
Table 1: Summary of pure models. These models employ the pure form of ViT, aiming to enhance its performance or fine-tune it. (Reported metrics are directly extracted from the original publications. In certain studies, per-class metrics are provided; metrics that are not reported are indicated with a "-".)
@@ 4.2 Hybrid Models
[41] Self-built 99.10% 0.9777 0.9847 0.9775 Using Multi-Head Self-Attention (MSA) and a Spatial Convolutional Self-Attention-based Transformer (SCSA-Transformer) for accurate and efficient recognition of multiple strawberry disease classes
@@ 5.1 Key Findings
  • Self-attention mechanism is a key advantage: The self-attention mechanism allows ViTs to focus on the most important parts of an image without manual feature extraction. This is helpful in detecting small, subtle spots on leaves. ViTs can dynamically weigh different regions of the input image, ensuring that significant features receive adequate attention during classification. This helps capture both local and global contextual information. This approach is different from CNNs, which apply convolution operations uniformly across an image and can sometimes miss small but significant features. The ability of ViTs to focus on key features helps improve accuracy and reduce the need for extensive pre-processing of images.
@@ 5.2 Open Challenges
Despite the significant progress made with ViTs in precision agriculture, several challenges still hinder their wider adoption. A key issue is the absence of inductive biases in ViTs. Unlike CNNs, which take advantage of spatial hierarchies and local patterns in images through convolutional filters, ViTs do not possess these built-in inductive biases. Consequently, ViTs require large, high-quality datasets to perform well, making them data-hungry. This reliance on vast amounts of data can be a major obstacle in agriculture, where acquiring labeled datasets is both costly and time-consuming. As a result, ViTs may struggle to reach optimal performance without substantial data augmentation or pre-training on large-scale datasets like ImageNet.
@@
  • Need for more comprehensive and diverse datasets: Many studies use publicly available datasets such as PlantVillage, which may not fully represent the variability of real-world scenarios. Real-world conditions involve a multitude of factors, such as diverse plant species, various disease types and stages, and different environmental and lighting conditions—factors that are often not captured in standard datasets. Datasets also need to account for image variations, including different angles, distances, and image quality. The lack of diverse data can lead to models that are not robust and may not generalize well to new, unseen data. Although some studies have created their own datasets to address this issue, collecting and annotating large-scale, high-quality datasets remains expensive and time-consuming. Table LABEL:tab:tabledata shows the number of times each dataset has been used.
@@
Table 3: Frequency of dataset usage across the analyzed studies. Datasets created by gathering images from the internet or by combining other datasets are categorized under the ’Custom’ category. Datasets created by the authors themselves are categorized under the ’Self-built’ category. Datasets that do not have a unique name or DOI link and are simply accessible from Kaggle are categorized under the ’From Kaggle’ category.
@@
Self-built Dataset built by the authors $12$
@@
Custom Datasets gathered from search engines, created by combining other datasets, or not easily accessible $8$
@@
CD&S [3] Contains over 4,000 corn images comprising field images and augmented images $2$
@@
DRLI [15] Contains over 4,000 images of healthy and diseased leaves $1$
PlantComposite [56] Contains over 58,000 images of healthy and diseased leaves $1$
Citrus Fruits and Leaves dataset [62] Contains over 600 images of citrus leaves and 150 images of citrus fruits, with five classes for each $1$
Dhan-Shomadhan [31] Contains over 1,100 images of five rice leaf diseases $1$
Strawberry Disease dataset [1] Contains 2,500 images of seven types of strawberry diseases $1$